OpenStudy (anonymous):

Number of natural numbers between 100 and 1000 such that at least one of their digits is 7, is?

OpenStudy (anonymous):

107 , 117 , ... 197 10 natural numbers 170 171 ... 179 10 natural numbers , but we counted '177' two times , so there are 19 natural number at least one of their digits is 7 between 100-200 It's the same with 100 , 200 , 300 ,...900 , except 700-800 8*19=152 and there are 100 ones between700-800, so answer is 252

OpenStudy (nikvist):

including 100 and 1000 or not?

OpenStudy (nikvist):

without 7: 8*9*9=648 total: 9*10*10=900 with 7: 900-648=252